zarasht:

how many hydrogen atoms are in 0.1488 g of phosphoric acid, H3PO4?

justjm:

Use dimensional analysis \(0.1488~g~H3PO4 * \frac{1~mol~H3PO4}{MM~H3PO4} * \frac{3~mol~H}{1~mol~H3PO4} * \frac{6.23E23~molecules}{1~mol~H} \)